# UniCATS: A Unified Context-Aware Text-to-Speech Framework with Contextual VQ-Diffusion and Vocoding

### v1

@@ Introduction
The typical neural text-to-speech (TTS) pipeline, such as Tacotron 2 (Shen et al. 2018) and FastSpeech 2 (Ren et al. 2020), consists of two stages: predicting the mel-spectrogram from text and then vocoding it into waveform. Additional techniques, such as normalizing flow (Valle et al. 2020; Kim et al. 2020) and diffusion models (Liu et al. 2022; Popov et al. 2021; Liu, Guo, and Yu 2023), have been introduced to generate the mel-spectrogram. Recently, VQTTS (Du et al. 2022) proposes a novel approach by utilizing discrete speech tokens as the intermediate representation for text-to-speech synthesis. The discrete tokens have been found to exhibit superior naturalness and robustness compared to mel-spectrograms. Textless NLP (Lakhotia et al. 2021) and AudioLM (Borsos et al. 2022) propose to leverage wav2vec 2.0 and w2v-BERT (Chung et al. 2021) respectively for language model training and consequently are able to generate speech unconditionally via auto-regressive inference. VALL-E (Wang et al. 2023) and SPEAR-TTS (Kharitonov et al. 2023) further extend the use of discrete tokens to zero-shot speaker adaptation. Specifically, they generate acoustic tokens based on the input text using a decoder-only auto-regressive (AR) model. During inference, they conduct AR continuation from the acoustic tokens of a short speech prompt provided by the target speaker. As a result, these models are capable of generating speech in the target speaker’s voice. NaturalSpeech 2 (Shen et al. 2023) employs a typical diffusion model to generate discrete acoustic tokens as continuous features.
@@
However, current TTS models that based on discrete speech tokens face three limitations. Firstly, most of these models are autoregressive (AR) models, which restricts them to generate speech only in a left-to-right direction. This limitation makes them unsuitable for speech editing, where both preceding and following contexts are provided. Secondly, the construction of acoustic tokens involves residual vector quantization (RVQ), resulting in multiple indices for each frame. This approach introduces prediction challenges and complexity. For instance, VALL-E incorporates a non-auto-regressive (NAR) module to generate the residual indices, while SPEAR-TTS addresses this issue by padding the RVQ indices into a longer sequence, which further complicates modeling. Lastly, the audio quality of these TTS systems is constrained by the performance of audio codec models.
@@
HifiGAN Semantic token X-vector 4.30 $\pm$ 0.08 3.96 $\pm$ 0.08 0.742
AudioLM Semantic token AR continuation 3.99 $\pm$ 0.07 3.96 $\pm$ 0.08 0.766
CTX-vec2wav Semantic token Contextual vocoding 4.75 $\pm$ 0.06 4.50 $\pm$ 0.07 0.818
@@
LibriTTS is a multi-speaker transcribed English speech dataset. Its training set consists of approximately 580 hours of speech data from 2,306 speakers. For evaluation purposes, we exclude 500 utterances from the official LibriTTS training set, which will serve as one of our test sets referred to as “test set A”. Test set A comprises 369 speakers out of the 2,306 training speakers. In addition, we utilize 500 utterances from the ”test-clean” set of LibriTTS, designated as ”test set B,” to assess the zero-shot adaptation capability for new and unseen speakers. Test set B contains 37 speakers. Each speaker in test sets A and B is associated with a brief speech prompt lasting approximately 3 seconds. The utterance list for both test sets A and B, along with their corresponding prompts, is available on our demo page. Lastly, for evaluating speech editing, we employ the same test set as utilized in (Yin et al. 2022). The utterances for this evaluation are also derived from the ”test-clean” set of LibriTTS and are denoted as ”test set C”.
@@
. CTX-txt2vec is trained for 50 epochs using an AdamW (Loshchilov and Hutter 2017) optimizer with a weight decay of $4.5\times 10^{-2}$. In CTX-vec2wav, both semantic encoders consist of $M=2$ Conformer-based blocks. The attention layers within these blocks have 2 heads and a dimension of 184. The mel encoder employs a 1D convolution with a kernel size of 5 and an output channel of 184. CTX-vec2wav is trained using an Adam (Kingma and Ba 2014) optimizer for 800k steps. The initial learning rate is set to $2\times 10^{-4}$ and is halved every 200k steps.
@@
Ground-truth 4.89 $\pm$ 0.04 4.54 $\pm$ 0.08 0.806 4.91 $\pm$ 0.04 4.50 $\pm$ 0.08 0.818
FastSpeech 2 3.81 $\pm$ 0.08 3.94 $\pm$ 0.06 0.775 3.65 $\pm$ 0.08 3.65 $\pm$ 0.07 0.775
VALL-E 4.23 $\pm$ 0.07 3.98 $\pm$ 0.06 0.794 4.17 $\pm$ 0.09 3.83 $\pm$ 0.07 0.792
UniCATS 4.54 $\pm$ 0.07 4.56 $\pm$ 0.07 0.799 4.43 $\pm$ 0.08 4.25 $\pm$ 0.08 0.806
@@
The results are shown in Table 1. Our proposed CTX-vec2wav demonstrates the best performance in speech resynthesis from semantic tokens in terms of both naturalness and speaker similarity. In contrast, when vocoding semantic tokens with HifiGAN, we observe the lowest SECS score. This can be attributed to the information compression inherent in x-vectors as a bottleneck feature of the speaker-verification model. Although x-vectors effectively distinguish between speakers, they are not ideal for accurately reconstructing the speaker’s voice. Remarkably, CTX-vec2wav even outperforms Encodec in subjective evaluations and achieves a SECS score comparable to the ground truth. We notice that Encodec resynthesis occasionally introduces artifacts that negatively impact the subjective scores. However, it still maintains a slightly higher SECS score than the ground truth. This phenomenon may be attributed to the robustness of the SECS evaluation model against artifacts.

### v3

@@ Introduction
The typical neural text-to-speech (TTS) pipeline, such as Tacotron 2 (Shen et al. 2018) and FastSpeech 2 (Ren et al. 2020), consists of two stages: predicting the mel-spectrogram from text and then vocoding it into waveform. Additional techniques, such as normalizing flow (Valle et al. 2020; Kim et al. 2020) and diffusion models (Liu et al. 2022; Popov et al. 2021; Liu, Guo, and Yu 2023), have been introduced to generate the mel-spectrogram. Recently, VQTTS (Du et al. 2022) proposes a novel approach by utilizing discrete speech tokens as the intermediate representation for text-to-speech synthesis. The discrete tokens have been found to exhibit superior naturalness and robustness compared to mel-spectrograms. Textless NLP (Lakhotia et al. 2021) and AudioLM (Borsos et al. 2022) propose to leverage wav2vec 2.0 and w2v-BERT (Chung et al. 2021) respectively for language model training and consequently are able to generate speech unconditionally via auto-regressive inference. InstructTTS (Yang et al. 2023) uses VQ-diffusion to generate acoustic tokens whose speaking style is guided by a natural language prompt. VALL-E (Wang et al. 2023) and SPEAR-TTS (Kharitonov et al. 2023) further extend the use of discrete tokens to zero-shot speaker adaptation. Specifically, they generate acoustic tokens based on the input text using a decoder-only auto-regressive (AR) model. During inference, they conduct AR continuation from the acoustic tokens of a short speech prompt provided by the target speaker. As a result, these models are capable of generating speech in the target speaker’s voice. NaturalSpeech 2 (Shen et al. 2023) employs a typical diffusion model to generate discrete acoustic tokens as continuous features.
@@
However, current TTS models that based on discrete speech tokens face three limitations. Firstly, most of these models are autoregressive (AR) models, which restricts them to generate speech only in a left-to-right direction. This limitation makes them unsuitable for speech editing, where both preceding and following contexts are provided. Secondly, the construction of acoustic tokens involves residual vector quantization (RVQ), resulting in multiple indices for each frame. This approach introduces prediction challenges and complexity into text-to-speech. For instance, VALL-E incorporates a non-auto-regressive (NAR) module to generate the residual indices, while SPEAR-TTS addresses this issue by padding the RVQ indices into a longer sequence, which further complicates modeling. Lastly, the audio quality of these TTS systems is constrained by the performance of audio codec models.
@@
HifiGAN Semantic token X-vector 4.30 $\pm$ 0.08 3.96 $\pm$ 0.08 0.764
AudioLM Semantic token AR continuation 3.99 $\pm$ 0.07 3.96 $\pm$ 0.08 0.787
CTX-vec2wav Semantic token Contextual vocoding 4.75 $\pm$ 0.06 4.50 $\pm$ 0.07 0.840
@@
LibriTTS is a multi-speaker transcribed English speech dataset. Its training set consists of approximately 580 hours of speech data from 2,306 speakers. For evaluation purposes, we exclude 500 utterances from the official LibriTTS training set, which will serve as one of our test sets referred to as “test set A”. Test set A comprises 369 speakers out of the 2,306 training speakers. In addition, we utilize 500 utterances from the “test-clean” set of LibriTTS, designated as “test set B”, to assess the zero-shot adaptation capability for new and unseen speakers. Test set B contains 37 speakers. Each speaker in test sets A and B is associated with a brief speech prompt lasting approximately 3 seconds. The utterance list for both test sets A and B, along with their corresponding prompts, is available on our demo page. Lastly, for evaluating speech editing, we employ the same test set as utilized in (Yin et al. 2022). The utterances for this evaluation are also derived from the “test-clean” set of LibriTTS and are denoted as “test set C”.
@@
. CTX-txt2vec is trained for 50 epochs using an AdamW (Loshchilov and Hutter 2017) optimizer with a weight decay of $4.5\times 10^{-2}$. The number of diffusion steps is set to $T=100$. In CTX-vec2wav, both semantic encoders consist of $M=2$ Conformer-based blocks. The attention layers within these blocks have 2 heads and a dimension of 184. The mel encoder employs a 1D convolution with a kernel size of 5 and an output channel of 184. CTX-vec2wav is trained using an Adam (Kingma and Ba 2014) optimizer for 800k steps. The initial learning rate is set to $2\times 10^{-4}$ and is halved every 200k steps.
@@
Ground-truth 4.89 $\pm$ 0.04 4.54 $\pm$ 0.08 0.832 4.91 $\pm$ 0.04 4.50 $\pm$ 0.08 0.843
FastSpeech 2 3.81 $\pm$ 0.08 3.94 $\pm$ 0.06 0.820 3.65 $\pm$ 0.08 3.65 $\pm$ 0.07 0.765
VALL-E 4.23 $\pm$ 0.07 3.98 $\pm$ 0.06 0.795 4.17 $\pm$ 0.09 3.83 $\pm$ 0.07 0.775
UniCATS 4.54 $\pm$ 0.07 4.56 $\pm$ 0.07 0.831 4.43 $\pm$ 0.08 4.25 $\pm$ 0.08 0.832
@@
The results are shown in Table 1. Our proposed CTX-vec2wav demonstrates the best performance in speech resynthesis from semantic tokens in terms of both naturalness and speaker similarity. In contrast, when vocoding semantic tokens with HifiGAN, we observe the lowest SECS score. This can be attributed to the information compression inherent in x-vectors as a bottleneck feature of the speaker-verification model. Although x-vectors effectively distinguish between speakers, they are not ideal for accurately reconstructing the speaker’s voice. Remarkably, CTX-vec2wav even outperforms Encodec in subjective evaluations and achieves a SECS score comparable to the ground truth. We notice that Encodec resynthesis occasionally introduces artifacts that negatively impact the subjective scores.